OpenStudy (anonymous):

How do i find the derivative of 3pi Sec^3[Pi x]Tan[Pi x]?

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ 3π\sec^3(πx)\tan(πx) }\) \(\LARGE\color{blue}{ 3π\frac{1}{\cos^3(πx)}\frac{\sin(πx)}{\cos(πx)} }\) \(\LARGE\color{blue}{ \frac{3π\sin(πx)}{\cos^4(πx)} }\) this is the f(x).

OpenStudy (solomonzelman):

(Refresh when you see weird things with question marks)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{3π\sin(πx)}{\cos^4(πx)} }\) take the derivative of \(\LARGE\color{blue}{ \frac{\sin(πx)}{\cos^4(πx)} }\) and multiply times 3pi. \(\LARGE\color{blue}{ \frac{\sin(πx)}{\cos^4(πx)} }\) d/dx is going to be: \(\LARGE\color{blue}{ \frac{cos^4(πx)~π\cos(πx)-4π\sin^3(πx)~\sin(πx)}{\cos^8(πx)} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{πcos^5(πx)-4π\sin^4(πx)}{\cos^8(πx)} }\)

OpenStudy (solomonzelman):

you have to REFRESH if you see wired rhombuses with question marks.